# boxparampadding=1.5pt, padding-left=2pt, padding-right=2pt, height=7.5pt, background-color=lightgray, border-radius=2pt, border-right-width=1pt, border-bottom-width=1pt, border-color=gray ExploreSelf: Fostering User-driven Exploration and Reflection on Personal Challenges with Adaptive Guidance by Large Language Models

### v2

@@ 7.3. Design Considerations for Long-term, Multi-session Interaction
As the first step of investigating the feasibility of leveraging LLM-driven exploration of personal challenges, our investigation was based on a lab study conducted in a single session. That being said, all participants except one envisioned using ExploreSelf continuously in the future, particularly when they needed to process complex thoughts or emotions. For example, P14 imagined the system could help users recognize recurring themes, providing deeper insights into long-standing thought patterns and behaviors. Their feedback underscores ExploreSelf’s potential to become a regular tool for self-reflection and processing personal challenges.
@@
However, applying ExploreSelf for long-term engagement—revisiting ExploreSelf in multiple sessions over time—involves new design challenges and opportunities. First, the system should incorporate sort of long-term memory (c.f., (Jo et al., 2024; Bae et al., 2022)), where it leverages the information derived from past sessions for the current session. For example, the system may memorize the interpersonal relationships of the user and consult it when generating themes and questions accordingly. However, simply stitching the entire interaction history for model input is neither scalable nor reliable (Bae et al., 2022). This calls for future research designing how the system constructs and synthesizes the user knowledge that are noteworthy.
@@
Second, prolonged user engagement with the system will also increase the volume of interaction history to be presented. Therefore, interaction techniques such as semantic zooming (Boulos, 2003) and the multi-level summarization approach would be necessary. The long-term data will also pose opportunities to present other supplementary data, such as recurring questions or themes that the user keeps avoiding exploring: P13 imagined, “I noticed that I have avoided the questions asking about a specific topic. Seeing this allowed me to better understand myself regarding what I don’t want to face.” Supporting such ‘byproduct’ reflections calls for further investigation from the perspectives of the personal informatics systems (Li et al., 2011; Choe et al., 2017; Kim et al., 2021b).
@@ 7.4. Limitations and Future Work
Considering the technical reliability and its availability, we selected OpenAI GPT—an LLM trained with datasets dominated by English and Western cultural references—as an underlying LLM for our generative pipelines in Korean. Although we did not encounter participants’ feedback that the AI generations did not correspond with Korean cultural sensitivity, using LLMs aligned with the language and cultural context of the study population (e.g., HyperCLOVA (Kim et al., 2021a) in Korean) may potentially improve the quality of generation (Seo et al., 2024; Naous et al., 2024).
@@ 8. Conclusion
We presented ExploreSelf, an LLM-driven application that supports processing personal challenges by exploring related themes and questions. ExploreSelf enables people to take the initiative of exploration, thereby maintaining their agency and proactively selecting what aspects they want to explore. To examine how people use ExploreSelf to explore and reflect on their own challenges, we conducted an exploratory user study with 19 individuals. Participants successfully adopted the concept of proactive exploration interactions that ExploreSelf offered, showing various patterns. Our study revealed that providing multiple options with adaptive guidance enhanced participants’ perceived agency, and the large language model played a key role by providing wordings that acknowledged the participants’ language in detail. Our work contributes to the growing body of LLM-driven mental well-being support, demonstrating the feasibility and usefulness of flexible user interface inspired by reflective writing.
@@ Acknowledgements.
We thank our participants in both formative interviews and the user study for their time and efforts. We are also grateful to Eunjoo Kim at Yonsei University and Kyungah Lee at Dodakim Child Development Center for helping us recruit professionals for our formative study. We also thank Dasom Choi for her feedback on our early version of draft. This work was supported as a research internship at NAVER AI Lab of NAVER Cloud.References
@@
- Heer (2019) Jeffrey Heer. 2019. Agency plus automation: Designing artificial intelligence into interactive systems. Proceedings of the National Academy of Sciences 116, 6 (2019), 1844–1850.
@@
- Hwang and Won (2022) Angel Hsing-Chi Hwang and Andrea Stevenson Won. 2022. Ai in your mind: Counterbalancing perceived agency and experience in human-ai interaction. In Chi conference on human factors in computing systems extended abstracts. 1–10.
@@
- Jain et al. (2018) Mohit Jain, Ramachandra Kota, Pratyush Kumar, and Shwetak N. Patel. 2018. Convey: Exploring the Use of a Context View for Chatbots. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–6. https://doi.org/10.1145/3173574.3174042
@@
- Lee et al. (2024) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L.C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Antonette Shibani, Disha Shrivastava, Lila Shroff, Agnia Sergeyuk, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia Ha Rim Rho, Zejiang Shen, and Pao Siangliulue. 2024. A Design Space for Intelligent and Interactive Writing Assistants. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 1054, 35 pages. https://doi.org/10.1145/3613904.3642697
@@
- Moruzzi and Margarido (2024) Caterina Moruzzi and Solange Margarido. 2024. Customizing the balance between user and system agency in human-AI co-creative processes. In Proceedings of the 15th International Conference on Computational Creativity (ICCC’24). Association for Computational Creativity. 1–10.
@@
- OpenAI (2024) OpenAI. 2024. OpenAI API. Retrieved Sep 10, 2024 from https://openai.com/api/
@@
Appendix A Summarized Task Instructions for Each Generation Pipeline
A.1. Theme Generation
@@
¡initial_information/¿: Client’s initial brief introductory of difficulty narrative, and the client’s background.
¡previous_session_log¿: Logs of sessions before the current session. DO NOT overlap with the previously selected themes.

### v3

@@ 7.2. Design Considerations for Long-term, Multi-session Interaction
At an early stage of investigating the feasibility of leveraging LLM-guided exploration of personal challenges, our work was based on a single-session lab study. That being said, in the debriefing, all participants except one expressed willingness to engage with ExploreSelf continuously in the future, particularly in circumstances where they would need to process complex thoughts or emotions. For example, P14 imagined the system could help users recognize recurring themes, providing deeper insights into long-standing thought patterns and behaviors. Their feedback underscores ExploreSelf’s potential to serve as a channel for people to engage with regularly to process their personal challenges.
@@
Supporting long-term engagement with ExploreSelf—revisiting ExploreSelf in multiple sessions over time—requires a new set of design considerations. The current version of ExploreSelf expands the branches of reflection rooted in the initial narrative, which the user provides in a prelude to the exploration. Intuitively treating this unit as a session, users may create multiple sessions or append more themes to a previously-explored session. Notably, the sessions of the same user may be interconnected, sharing common aspects of and sometimes influencing one another. For example, a series of personal challenges can lead to the discovery of an overarching theme. Also to ensure accuracy and reliability, the system should incorporate sort of long-term memory (c.f., (Jo et al., 2024; Bae et al., 2022)), where it leverages the information derived from past sessions for the current session. For example, the system may memorize the interpersonal relationships of the user and consult it when generating themes and questions accordingly. However, simply stitching the entire interaction history for model input is neither scalable nor reliable (Bae et al., 2022), and could lead to inaccurate or irrelevant summaries and guidance. This calls for future research in designing how the system constructs and synthesizes the user knowledge that is noteworthy.
@@
Also, prolonged user engagement with the system will also increase the volume of interaction history to be presented. Therefore, interaction techniques such as semantic zooming (Boulos, 2003) and the multi-level summarization approach would be necessary. The long-term data will also pose opportunities to present other supplementary data, such as recurring questions or themes that the user keeps avoiding exploring: P13 imagined, “I noticed that I have avoided the questions on a specific topic. Seeing this allowed me to better understand myself about what I don’t want to face.” Supporting such ‘byproduct’ reflections calls for further investigation from the perspectives of the personal informatics systems (Li et al., 2011; Choe et al., 2017; Kim et al., 2021b).
To comprehensively support people in addressing personal challenges, future versions of ExploreSelf should go beyond the exploration phase. Our user study results indicate that about 30 minutes of exploration fostered a certain level of sensemaking about the self and the subject matter. According to the transtheoretical model (Prochaska and Velicer, 1997), the exploration mostly supports individuals to reach the contemplation stage, where they recognize what to change to address their challenge. However, different types of guidance and interaction may be more effective for those who have progressed to the preparation stage, where they began taking small steps towards change.
7.3. Role and Stance of AI in Guiding Reflection
Although we minimized conversational affordances from the interface, ExploreSelf still conveyed messages in human language with a tone through comments. Within the current scope of investigation, we followed a broadly applicable guideline and instructed the model to generate empathetic and supportive comments. However, the positive tone of comments was sometimes perceived as inconsiderate of the gravity of the participants’ concerns. This underscores the need for a more nuanced design of the AI’s role and stance embedded in its messages. While ExploreSelf is not intended to mimic psychotherapy, the therapists from our formative interviews highlighted that balancing both empathy and confrontation is often a helpful therapeutic tactic, pointing out that they do not always show agreement and support to their clients in leading them to positive change. Notably, counselors sometimes confront their clients with discrepancies in their perception of reality (Anderson, 1968), which may lead to resistance (Resnicow and McMaster, 2012).
Applied to ExploreSelf, the AI may take a stance that properly facilitates reflection on the fly that is aligned with these principles. To enable users to engage with confronting guidance when they are mentally ready, the system may provide a set of different stances that users can select as needed. However, LLMs, particularly those fine-tuned with human alignment techniques like RLHF (Ouyang et al., 2022), typically exhibit supportive, empathetic, and cooperative behaviors (Wei et al., 2024; Lee et al., 2024b; Zhu et al., 2024) and this tendency is difficult to override by prompting alone (Li et al., 2024; Zheng et al., 2024). To demonstrate a confronting stance, the system may leverage customized LLMs fine-tuned with annotated transcripts from psychotherapy sessions such as Motivational Interviewing (Resnicow and McMaster, 2012). Here, users’ interaction logs would also play a key role in assessing their resistance to specific themes. For instance, P12 and P14 noted that reviewing unselected questions might uncover the aspects of the challenge that he often avoids contemplating. A future version of ExploreSelf can analyze the patterns in these roads not taken and determine the level of confrontational guidance provided to the user.
7.4. Supporting Severe and Sensitive Challenges
Observing participants exploring diverse personal challenges of varying severity, we identified implications for supporting individuals with severe challenges, including those related to mental illness, such as depression, traumas, or other emotional distress. For individuals experiencing severe apathy, commonly seen in conditions like depression, self-directed exploration may be overwhelming (Smith et al., 2022; Kim et al., 2024a). This echoes the feedback from experts in our formative interviews, who noted that individuals with severe emotional numbness may struggle to engage meaningfully with open-ended reflective tools. In such cases, more proactive guidance can be more effective than a high degree of choice.
Similarly, reflecting on one’s trauma necessitates careful guidance due to its emotional intensity involved (Perlman, 1998; Miller, 2001). To keep the exploration within the emotionally safe boundary, ExploreSelf may receive traumatic keywords or precautions from users in advance to consider them when handling the emotionally charged material. Although ExploreSelf allows users to decide when to stop drilling into certain themes, individuals exploring a traumatic challenge may become overly involved in the exploration. Future work can investigate whether and how the system should intervene during exploration. However, simply terminating a session, particularly when users lack alternative resources, raises a lot of ethical concerns about potentially leaving individuals without support in critical moments. Instead, it is essential to develop guardrails that facilitate navigation for vulnerable users while ensuring their safety. This is especially important given that many individuals increasingly turn to accessible LLM interfaces for coping with mental health challenges in the absence of other available options (Song et al., 2024).
Cultural awareness is crucial for addressing deeply complex and sensitive challenges, as cultural consideration is an integral aspect of mental health support (Pendse et al., 2019; Balcombe and De Leo, 2022; Christopher, 1999; Lent, 2004). Considering the technical reliability and availability, we selected OpenAI GPT—an LLM trained with datasets dominated by English and Western cultural references—as the underlying model for our generative pipelines in Korean. Although we did not receive participants’ feedback that the AI generations lacked sensitivity to Korean cultural norms, the system may still reflect certain Western cultural norms and biases (Pawar et al., 2024; Tao et al., 2024) when handling challenges specific to non-Western contexts. This limitation can be particularly problematic when engaging with marginalized communities, for whom issues such as racism or sexism are not mere distortions but ongoing harms (Haslanger, 2017; McGowan, 2019). Employing LLMs aligned with the language and cultural context of the target population (e.g., HyperCLOVA (Kim et al., 2021a) in Korean) could better reflect culturally relevant references (Seo et al., 2024; Naous et al., 2024). However, the disparity in the model performance of LLMs across different languages and regions remains an open challenge.
@@ 8. Conclusion
We presented ExploreSelf, an LLM-driven application that supports processing personal challenges by exploring related themes and questions. ExploreSelf enables people to take the initiative of exploration, thereby maintaining their agency and proactively selecting what aspects they want to explore. To examine how people use ExploreSelf to explore and reflect on their own challenges, we conducted an exploratory user study with 19 individuals. Participants successfully adopted the concept of proactive exploration interactions that ExploreSelf offered, showing various patterns. Our study revealed that providing multiple options with adaptive guidance enhanced participants’ perceived agency, and the large language model played a key role by providing wordings that acknowledged the participants’ language in detail. We highlighted several areas for future research, including designing LLM-guided self-reflection systems that balance autonomy and intervention, technical and design considerations for long-term reflective practices, and incorporating cultural and emotional sensitivity to address diverse user needs and challenges. In summary, our work contributes to the growing body of LLM-driven mental well-being support, demonstrating the feasibility and usefulness of flexible user interface inspired by reflective writing.
@@ Acknowledgements.
We thank our participants in both formative interviews and the user study for their time and efforts. We are also grateful to Eunjoo Kim at Yonsei University and Kyungah Lee at Dodakim Child Development Center for helping us recruit professionals for our formative study. We also thank Dasom Choi for her feedback on the early version of our draft. This work was supported through a research internship at NAVER AI Lab of NAVER Cloud.References
@@
- Anderson (1968) Susan C Anderson. 1968. Effects of Confrontation by High- and Low-functioning Therapists. Journal of Counseling Psychology 15, 5p1 (1968), 411.
@@
- Balcombe and De Leo (2022) Luke Balcombe and Diego De Leo. 2022. Human-computer interaction in digital mental health. In Informatics, Vol. 9. MDPI, 14.
- Bandura (1991) Albert Bandura. 1991. Social cognitive theory of self-regulation. Organizational behavior and human decision processes 50, 2 (1991), 248–287.
@@
- Christopher (1999) John Chambers Christopher. 1999. Situating psychological well-being: Exploring the cultural roots of its theory and research. Journal of Counseling & Development 77, 2 (1999), 141–152.
@@
- Cleary and Zimmerman (2004) Timothy J Cleary and Barry J Zimmerman. 2004. Self-regulation empowerment program: A school-based program to enhance self-regulated and self-motivated cycles of student learning. Psychology in the Schools 41, 5 (2004), 537–550.
@@
- Deci and Ryan (2013) Edward L Deci and Richard M Ryan. 2013. Intrinsic motivation and self-determination in human behavior. Springer Science & Business Media.
@@
- Google, Inc (2023) Google, Inc. 2023. Bard - Chat Based AI Tool from Google, Powered by PaLM 2. Retrieved Sep 10, 2024 from https://bard.google.com/
@@
- Haslanger (2017) Sally Haslanger. 2017. Racism, ideology, and social movements. (2017).
- Hayes and Pierson (2005) Steven C Hayes and Heather Pierson. 2005. Acceptance and Commitment Therapy. Springer.
@@
- Lee et al. (2024a) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L.C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Antonette Shibani, Disha Shrivastava, Lila Shroff, Agnia Sergeyuk, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia Ha Rim Rho, Zejiang Shen, and Pao Siangliulue. 2024a. A Design Space for Intelligent and Interactive Writing Assistants. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 1054, 35 pages. https://doi.org/10.1145/3613904.3642697
@@
- Lee et al. (2024b) Yoon Kyung Lee, Jina Suh, Hongli Zhan, Junyi Jessy Li, and Desmond C. Ong. 2024b. Large Language Models Produce Responses Perceived to be Empathic. arXiv:2403.18148 [cs.CL] https://arxiv.org/abs/2403.18148
- Lent (2004) Robert W Lent. 2004. Toward a unifying theoretical and practical perspective on well-being and psychosocial adjustment. Journal of counseling psychology 51, 4 (2004), 482.
@@
- Li et al. (2024) Wenkai Li, Jiarui Liu, Andy Liu, Xuhui Zhou, Mona Diab, and Maarten Sap. 2024. BIG5-CHAT: Shaping LLM Personalities Through Training on Human-Grounded Data. arXiv:2410.16491 [cs.CL] https://arxiv.org/abs/2410.16491
@@
- McGowan (2019) Mary Kathryn McGowan. 2019. Just words: on speech and hidden harm. Oxford University Press.
@@
- Miller (2001) Madelyn Miller. 2001. Creating a safe frame for learning: Teaching about trauma and trauma treatment. Journal of Teaching in Social Work 21, 3-4 (2001), 159–176.
@@
- OpenAI (2024a) OpenAI. 2024a. ChatGPT: Optimizing Language Models for Dialogue. Retrieved Sep 10, 2024 from https://openai.com/blog/chatgpt/
- OpenAI (2024b) OpenAI. 2024b. OpenAI API. Retrieved Sep 10, 2024 from https://openai.com/api/
@@
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback. arXiv:2203.02155 [cs.CL] https://arxiv.org/abs/2203.02155
@@
- Pawar et al. (2024) Siddhesh Pawar, Junyeong Park, Jiho Jin, Arnav Arora, Junho Myung, Srishti Yadav, Faiz Ghifari Haznitrama, Inhwa Song, Alice Oh, and Isabelle Augenstein. 2024. Survey of Cultural Awareness in Language Models: Text and Beyond. arXiv preprint arXiv:2411.00860 (2024).
- Pendse et al. (2019) Sachin R Pendse, Naveena Karusala, Divya Siddarth, Pattie Gonsalves, Seema Mehrotra, John A Naslund, Mamta Sood, Neha Kumar, and Amit Sharma. 2019. Mental health in the global south: challenges and opportunities in HCI for development. In Proceedings of the 2nd ACM SIGCAS Conference on Computing and Sustainable Societies. 22–36.
@@
- Perlman (1998) Stuart D Perlman. 1998. The therapist’s emotional survival: Dealing with the pain of exploring trauma. Jason Aronson, Incorporated.
@@
- Prochaska and Velicer (1997) James O. Prochaska and Wayne F. Velicer. 1997. The Transtheoretical Model of Health Behavior Change. American Journal of Health Promotion 12, 1 (1997), 38–48. https://doi.org/10.4278/0890-1171-12.1.38 arXiv:https://doi.org/10.4278/0890-1171-12.1.38 PMID: 10170434.
@@
- Resnicow and McMaster (2012) Ken Resnicow and Fiona McMaster. 2012. Motivational Interviewing: Moving from Why to How with Autonomy Support. International Journal of Behavioral Nutrition and Physical Activity 9 (2012), 1–9.
@@
- Ryan and Deci (2008) Richard M Ryan and Edward L Deci. 2008. A self-determination theory approach to psychotherapy: The motivational basis for effective change. Canadian psychology/Psychologie canadienne 49, 3 (2008), 186.
- Ryan et al. (2011) Richard M. Ryan, Martin F. Lynch, Maarten Vansteenkiste, and Edward L. Deci. 2011. Motivation and Autonomy in Counseling, Psychotherapy, and Behavior Change: A Look at Theory and Practice. The Counseling Psychologist 39, 2 (2011), 193–260. https://doi.org/10.1177/0011000009359313 arXiv:https://doi.org/10.1177/0011000009359313
- Scheel (2011) Michael J. Scheel. 2011. Client Common Factors Represented by Client Motivation and Autonomy. The Counseling Psychologist 39, 2 (2011), 276–285. https://doi.org/10.1177/0011000010375309 arXiv:https://doi.org/10.1177/0011000010375309
@@
- Smith et al. (2022) Ryan Smith, Samuel Taylor, Robert C Wilson, Anne E Chuning, Michelle R Persich, Siyu Wang, and William DS Killgore. 2022. Lower levels of directed exploration and reflective thinking are associated with greater anxiety and depression. Frontiers in Psychiatry 12 (2022), 782136.
@@
- Tao et al. (2024) Yan Tao, Olga Viberg, Ryan S Baker, and René F Kizilcec. 2024. Cultural bias and cultural alignment of large language models. PNAS nexus 3, 9 (2024), pgae346.
@@
- Zheng et al. (2024) Mingqian Zheng, Jiaxin Pei, Lajanugen Logeswaran, Moontae Lee, and David Jurgens. 2024. When ”A Helpful Assistant” Is Not Really Helpful: Personas in System Prompts Do Not Improve Performances of Large Language Models. arXiv:2311.10054 [cs.CL] https://arxiv.org/abs/2311.10054
- Zhu et al. (2024) Minjun Zhu, Linyi Yang, and Yue Zhang. 2024. Personality Alignment of Large Language Models. arXiv:2408.11779 [cs.CL] https://arxiv.org/abs/2408.11779
Appendix A Formative Study Material
A.1. Example Narrative
Ever since middle school, I’ve always seen myself as someone who thrives in the shadows, away from the glaring spotlight of public attention. I preferred to keep my head down, diligently working in quiet solitude, and my social interactions were always subtle, never the center of attention. But recently, my usual tranquility at work was disrupted when I was unexpectedly chosen to lead a major project. This wasn’t just any project; it involved a division-wide presentation, thrusting me into the very spotlight I had always avoided. The thought of standing in front of all those people, having to lead and direct, filled me with an intense dread that was completely new to me.
Appendix B Summarized Task Instructions for Each Generation Pipeline
B.1. Theme Generation
@@
<initial_information/>: Client’s initial brief introductory of difficulty narrative, and the client’s background.
<previous_session_log>: Logs of sessions before the current session. DO NOT overlap with the previously selected themes.